Suppose you borrow $600 of principal that must be repaid at the end of two years, alond with interest of 4 percent a year. If the annual inflation rate turns out to be 8 percent, (a) What is the real rate of interest on the loan? Instructions: Enter your response rounded to the nearest whole number. If you are entering any negative numbers be sure to include a negative sign () in front of those numbers percent (b) What is the real value of the principal repayment? Instructions: Enter your response rounded to two decimal places. (c) Who loses, the debtor or the creditor? Debtor Creditor Neither
Answer #1

Real interest on loan = (1.04/1.08)-1=-3.7% or -4%

Real value of principal repayment = 600/(1.08^2)=$514.40

Hence the creditor loses as the principal payment is less.